Hey guyz badly need help.

A parking meter slots receives dimes and nickels.when the meter box was emptied there were 148 coins and a total of $10.65.how many dimes and how many nickels were there.
please help.
thanks :)

Answer by mananth(16946)

x= number of dimes
y= number of nickels

1 x + 1 y = 180 .............1

10 x + 5 y = 1065 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-5 x -5 y = -900
10 x + 5 y = 1065
Add the two equations
5 x = 165
/ 5
x = 33
plug value of x in (1)
1 x + 1 y = 180
33 + y = 180
y = 180 -33
y = 147
y = 147
x= 33 number of dimes
y= 147 number of nickels
m.ananth@hotmail.ca